# Keldysh field theory of dynamical exciton condensation transitions in nonequilibrium electron-hole bilayers

### v2

@@ Abstract
If tunneling between layers is negligible, the potential difference required to maintain a nonzero steady state exciton density can be gauged away, so the system is equivalent to an equilibrium electron-hole bilayer with a reduced effective band gap. On the other hand, nonzero interlayer tunneling introduces a tunneling current [24, 38] that drives the system out of equilibrium, leading to new physics [40]. In this Letter we present a microscopic theory of the nonequilibrium exciton condensate based on a nonequilibrium field theory derived using the Keldysh formalism [41, 42, 43] that includes the effects of both a bias voltage and interlayer tunneling.
Depending on the symmetries of the bilayer crystal and the orbital character of each layer, interlayer tunneling may transform trivially (s-wave) or nontrivially (p-wave, d-wave, etc.) under rotations, or may be a spatially varying function [44]. We focus on the case in which interlayer tunneling is uniform in space, and assume p-wave interlayer tunneling for most of the explicit calculations in this Letter, with other forms of interlayer tunneling briefly discussed at the end. Different from s-wave tunneling, p-wave interlayer tunneling produces a potential landscape that is second order in the phase angle of the exciton field, leading when no bias voltage is applied to a second-order Josephson effect that breaks the interlayer phase symmetry down to $\mathbb{Z}_{2}$ from $U(1)$ [45]. We find that in the large bias limit, the system is described by an effective action in which interlayer $U(1)$ phase symmetry is effectively restored. We find that instead of Josephson effect, p-wave interlayer tunneling leads to a reduction of effective band gap and an increase of effective temperature for intervalley excitons.
@@
$H_{0}=\sum_{\tau k}\left(\frac{k^{2}}{2m}+\frac{E_{g}}{2}\right)(a_{\tau ck}^{\dagger}a_{\tau ck}-a_{\tau vk}^{\dagger}a_{\tau vk})$ (1)
describes the kinetic energy of conduction band electrons and valence band holes. Here $E_{g}$ is the band gap that can be tuned by a perpendicular electric field produced by the difference between top and bottom gate voltages, $\tau=\pm$ is the valley index, and $m$ is the effective masses of electrons and holes which are assumed identical for simplicity. The next term in the Hamiltonian
@@
describes interlayer tunneling. The momentum and valley dependence of $t_{\tau k}$ depends on symmetries of the system and is crucial for our upcoming results. We assume for most of the discussion below that interlayer tunneling is p-wave:
$t_{\tau k}=v_{t}(\tau k_{x}+ik_{y}),$ (3)
which is the case for InAs/GaSb quantum wells [46, 47, 48] and aligned TMD homobilayers with certain high-symmetry stacking registries [44, 49]. The Coulomb interaction term
$H_{C}=\frac{1}{2A}\sum_{bb^{\prime}\tau\tau^{\prime}}\sum_{kk^{\prime}q}V_{bb^{\prime}}(q)a_{\tau b,k+q}^{\dagger}a_{\tau^{\prime}b^{\prime},k^{\prime}-q}^{\dagger}a_{\tau^{\prime}b^{\prime}k^{\prime}}a_{\tau bk}$ (4)
distinguishes intralayer ($b=b^{\prime}$) and interlayer ($b\neq b^{\prime}$) interactions but neglects intervalley scattering due to the large momentum transfer required. Here $b,b^{\prime}=c,v$ are the band (layer) labels, and $A$ is the system area.
@@
Keldysh action.— We derive a nonequilibrium field theory that describes a biased electron-hole bilayer with interlayer tunneling based on the Keldysh formalism [41, 42, 43], outlining here the procedure to obtain the Keldysh action and presenting the main results, with detailed derivations left for the Supplemental Material. We express the model as a path integral along a closed time path $C$ that starts from the distant past, proceeds to the distant future, and then returns to the starting point. The generating function is
@@
where $\mathcal{T}_{C}$ is the contour ordering operator along $C$, and $\rho_{0}$ is the density matrix of the system in the distant past which we take as the equilibrium distribution of decoupled electron and hole layers: $\rho_{0}=e^{-(H_{0}-\mu_{c}N_{c}-\mu_{v}N_{v})/T}$ where $N_{b}=\sum_{\tau k}a_{\tau bk}^{\dagger}a_{\tau bk}$ is the number of electrons in each layer. For notational convenience Eq. (5) is written for a closed system; coupling to leads is included in the theory as the imaginary (dissipative) part of inverse Green functions as detailed in the Supplemental Material. To derive a theory of excitons, we perform a Hubbard-Stratonovich transformation of interlayer electron-hole interactions and introduce the electron-hole pairing fields $\Delta_{kq}^{\tau\tau^{\prime}}$, where $k$ and $q$ are respectively the relative momentum and center-of-mass momentum of an electron-hole pair, and $\tau,\tau^{\prime}$ are the valley indices of electrons and holes. The $k$-dependence of the pairing fields is irrelevant to the low-energy physics we discuss and is eliminated by projecting the $\Delta$ fields onto the 1s-exciton basis by defining
@@
$\frac{k^{2}}{m}\varphi_{k}-\frac{1}{A}\sum_{k^{\prime}}V_{cv}(k-k^{\prime})\varphi_{k^{\prime}}=-E_{b}\varphi_{k}.$ (7)
Here the exciton binding energy $E_{b}$ is defined as the absolute value of the 1s-exciton energy. The 1s-exciton fields $\Phi$ have two valley indices, one for electrons and the other for holes, and we express them in terms of a four-component spinor $(\Phi^{\mu})$ defined as $\Phi^{\tau\tau^{\prime}}=(\sum_{\mu}\Phi^{\mu}\tau_{\mu}/\sqrt{2})^{\tau\tau^{\prime}}$ where $\tau_{0}$ and $\tau_{1,2,3}$ are the $2\times 2$ identity and Pauli matrices in valley space [15]. In this notation $\Phi^{0},\Phi^{3}$ are intravalley exciton fields and $\Phi^{1},\Phi^{2}$ are intervalley exciton fields. Integrating out the fermion fields, we obtain an effective action in terms of the 1s-exciton fields $\Phi$. It is convenient to transform the forward ($+$) and backward (--) branches of the $\Phi$ fields into classical (c) and quantum (q) fields defined as
@@
$S_{0}[\bar{\Phi},\Phi]=\frac{1}{A}\sum_{q}\int\frac{d\omega}{2\pi}\,{\rm Tr}\bigg{[}(\omega-\frac{q^{2}}{4m}-E_{g}+E_{b}+i\gamma)\Phi^{\rm q\dagger}_{q}(\omega)\Phi^{\rm c}_{q}(\omega)+{\rm c.c.}+ig(\omega-\mu_{x})\coth\frac{\omega-\mu_{x}}{2T}\Phi^{\rm q\dagger}_{q}(\omega)\Phi^{\rm q}_{q}(\omega)\bigg{]}$ (10)
which describes free excitons with energy $E_{g}-E_{b}+q^{2}/4m$ and chemical potential $\mu_{x}$ at temperature $T$. While the quadratic coefficients take the stated form only in the dilute exciton regime (BEC regime) $|E_{g}-E_{b}-\mu_{x}|\ll E_{b}$ and in the frequency range $|\omega-\mu_{x}|\ll E_{b}$, the overall form of Eq. (10) is general and we expect that our qualitative results apply to all parameter regimes. The imaginary coefficients $\gamma$ and $g$ describe coupling of excitons to leads. Fluctuation-dissipation theorem implies the relation $\gamma=g(\omega-\mu_{x})$. In the absence of interlayer tunneling, the bias voltage $\mu_{x}$ can be absorbed into the dummy variable $\omega$ and the system is equivalent to an unbiased bilayer with a reduced band gap $E_{g}-E_{b}-\mu_{x}$. Excitons spontaneously form and undergo BEC at low enough temperatures when $E_{g}-E_{b}<\mu_{x}$. Below the transition the exciton fields have semiclassical solutions of the form $\Phi^{\rm c}(t)=|\Phi^{\rm c}|e^{-i\mu_{x}t}$ with amplitude determined by the ratio of quadratic and quartic coefficients of the action.
@@
in which the intervalley exciton fields $\Phi^{1},\Phi^{2}$ at frequency $\omega$ are coupled to those at frequency $-\omega$. Because of this coupling, the bias voltage $\mu_{x}$ cannot be absorbed into $\omega$ and the system gets out of equilibrium. Violation of fluctuation-dissipation theorem when $\mu_{x}\neq 0$ and $c_{J}\neq 0$ is explicitly shown in the Supplemental Material. Diagrammatic representations of $S_{0}$ and $S_{J}$ are shown in Fig. 2(a,b).
@@
Figure 2: Diagrammatic representations of the effective exciton action (a,b) and an electron-hole scattering process due to interlayer tunneling (c). The solid and dotted curves represent fermion and boson fields respectively, the crosses represent interlayer tunneling, and the wavy line represents Coulomb interaction. (a) and (b) respectively represent the free exciton action $S_{0}$ and Josephson action $S_{J}$. In the Keldysh formalism, the vertices and fermion lines are $2\times 2$ matrices in Keldysh space; more detailed diagrams are shown in the Supplemental Material.
If interlayer tunneling is s-wave, a first-order Josephson term proportional to $\Phi^{\rm q}(\omega=0)$ exists. Second-order terms of the form $\Phi(-\omega)\Phi(\omega)$ also exist and the coefficients are equal for all valley components. For p-wave interlayer tunneling (3), angular momentum conservation implies that first-order Josephson terms vanish and that second-order terms are nonzero only for intervalley exciton fields $\Phi^{1},\Phi^{2}$. Below we assume p-wave interlayer tunneling and leave the discussion of s-wave tunneling to the end of this Letter.
@@
For p-wave tunnelling, the Josephson action (11) produces an energy landscape with explicit dependence on the phase angle $\theta={\rm arg}\,\Phi$ of the form $E_{J}\sim\cos 2\theta$, the $U(1)$ phase symmetry is broken down to $\mathbb{Z}_{2}$ and the interlayer tunneling current satisfies the second-order Josephson relation $I\sim\sin 2\theta$ [45]. For an unbiased electron-hole bilayer below the BEC transition, the exciton fields are static and the system picks one of the two preferred phase angles that differ by $\pi$ as the ground state implying Ising-type phase transitions of the exciton fields. Because the Josephson action (11) involves only intervalley exciton fields, intervalley excitons are energetically favored over intravalley excitons by pinning the phase at one of the two preferred phase angles, in agreement with mean-field theory results in the context of InAs/GaSb quantum wells [46, 47].
@@
Large bias limit.— In the absence of interlayer tunneling, the phase of the exciton field rotates at a constant frequency $\omega=\mu_{x}$. Interlayer tunneling leads to a potential landscape that explicitly breaks the $U(1)$ phase symmetry. For small bias voltage $\mu_{x}$, the exciton condensate is a static one with its phase trapped at one of the potential minima. Above a threshold bias voltage $\mu_{x}\sim c_{J}$ the condensate becomes a dynamical one with rotating phase. The transition from static to dynamical condensates is schematically shown in Fig. 3. If the bias voltage is much larger than the Josephson energy scale $c_{J}$, the phase-dependent energy landscape is swept rapidly by the rotating fields at approximately constant frequency $\omega\approx\mu_{x}$. Instead of Josephson effects, the Josephson action produces an average effect on the exciton fields and $U(1)$ symmetry is effectively restored.
@@
Figure 3: Schematic phase diagram of a biased electron-hole bilayer. $E_{g}$ is the interlayer band gap that can be tuned by a displacement field, $E_{b}$ is the binding energy of interlayer excitons, and $\mu_{x}$ is the bias voltage. The blue region represents the region in which exciton condensation occurs, and the color scale represents the strength of excitonic coherence. The condensate is in the BEC regime at low exciton density and BCS regime at high density. In the presence of interlayer tunneling, below a threshold bias voltage the condensate is static with a constant phase, and above the threshold the condensate is dynamical with a rotating phase. The hatched area represents the large-bias BEC regime in which our theory applies.
@@
Figure 4: Exciton density $n_{x}$ as a function of bias voltage (exciton chemical potential) $\mu_{x}$. The solid and dashed curves are the $n_{x}$$\mu_{x}$ curves with and without interlayer tunneling respectively and are shifted horizontally by $\varepsilon$. $\mu_{0}=E_{g}-E_{b}$ is the threshold bias voltage for injection of excitons in the absence of interlayer tunneling, $\mu_{c}$ and $n_{c}$ are the critical bias voltage and critical density for the occurrence of BKT transition, while $\mu_{0}^{\prime}=\mu_{0}-\delta\mu_{0}$, $\mu_{c}^{\prime}=\mu_{c}-\delta\mu_{c}$, and $n_{c}^{\prime}=n_{c}+\delta n_{c}$ (see Eq. 14) are the corresponding quantities in the presence of p-wave interlayer tunneling.
A finer look at the nonequilibrium effects unveils that, despite the small $E_{b}/E_{g}$ ratio, the nonequilibrium dissipative term $\delta T$ can be as important as the $\varepsilon$ term. To see this, we sketch in Fig. 4 the density of intervalley excitons $n_{x}$ as a function of the bias voltage $\mu_{x}$. The gap reduction for intervalley exciton discussed above shifts the $n_{x}$$\mu_{x}$ curve to the left by $\delta\mu_{0}=\varepsilon$. Exciton condensation occurs when the temperature is below the Berezinskii-Kosterlitz-Thouless (BKT) transition temperature [55, 56, 14]
$T_{\rm BKT}\approx 1.3\,\frac{n_{x}}{m_{x}},$ (13)
with $m_{x}=2m$ the exciton mass and $n_{x}$ the exciton density. In other words, the critical exciton density for the occurrence of BKT transition at temperature $T$ is $n_{c}\approx m_{x}T/1.3$. Due to the effective temperature increase, the critical density increases by $\delta n_{c}\approx m_{x}\delta T/1.3$. Since the $n_{x}$$\mu_{x}$ curve is approximately linear at small exciton densities, with the slope approximately given by [15, 38] the geometric capacitance $C=e^{2}\partial n_{x}/\partial\mu_{x}\approx\epsilon/d$ of the bilayer, the critical bias voltage in the presence of interlayer tunneling decreases by
$\delta\mu_{c}=\delta\mu_{0}-\frac{e^{2}}{C}\delta n_{c}\approx\varepsilon-\frac{m_{x}e^{2}d}{1.3\,\epsilon}\delta T,$ (14)
@@
The assumption of p-wave interlayer tunneling (3) is crucial for our results and deserves further discussion. For any form of interlayer tunneling that transforms nontrivially (p-wave, d-wave, etc.) under rotation, the first-order Josephson terms vanish due to angular momentum conservation and the second-order terms are nonzero only when the two tunneling terms carry opposite chiralities, i.e., for intervalley excitons only. Our results therefore qualitatively apply for all higher-angular-momentum interlayer tunneling. On the other hand, when interlayer tunneling is s-wave, a nonzero first-order Josephson term ($\propto\Phi(\omega=0)$) exists for intravalley excitons, leading to nonzero static exciton density even before the condensation transition occurs. While the tunneling-induced static excitons are not coupled to the high-frequency exciton fields at quadratic level, electrostatic repulsion between excitons leads to an effective gap increase, i.e., a higher critical bias voltage for the onset of BEC.
Our theory applies to InAs/GaSb quantum wells in which excitons are formed at the $\Gamma$ valley and interlayer tunneling is p-wave due to different rotational symmetries of the conduction and valence band orbitals. For $K$-valley excitons in angle-aligned (R-type stacking) TMD homobilayers, the functional form of interlayer tunneling depends on the stacking registry [44, 57, 58]. For two of the three high-symmetry stackings ($R_{h}^{h}$ and $R_{h}^{X}$ [58]), interlayer tunneling is p-wave and our theory is directly applicable. For TMD heterobilayers or homobilayers with a nonzero twist angle, the two layers form a moiré pattern with spatially varying local stacking registry. A proper treatment of general TMD bilayers needs to take account of the momentum shift between conduction and valence bands [59, 44, 60, 49] and is left for future work. Excitonic coherence between shifted bands leads to density wave states that break translational symmetry [47, 61, 62]. In a simple intuitive picture, excitons in a moiré potential are localized near one of the high-symmetry stacking sites [63, 49], and the effects of interlayer tunneling are determined by the local stacking registry.

### v3

@@ Abstract
If tunneling between layers is negligible, the potential difference required to maintain a nonzero steady state exciton density can be gauged away, so the system is equivalent to an equilibrium electron-hole bilayer with a reduced effective band gap. Nonzero interlayer tunneling introduces a tunneling current [24, 38] that drives the system out of equilibrium, leading to new physics [40] different from that of driven-dissipative condensates [41, 42, 43, 44, 45]. In this Letter we present a microscopic theory of the nonequilibrium exciton condensate based on a Keldysh nonequilibrium field theory [46, 47, 48, 42] that includes the effects of both a bias voltage and interlayer tunneling.
In TMD bilayers, because the conduction and valence band extrema are located at the ±$\pm K$-valleys, the functional form of interlayer tunneling depends on the local stacking registry [49, 50, 51, 52]. In this Letter we focus on the experimentally relevant case of angle-aligned TMD homobilayers in which interlayer tunneling is uniform in space, and assume in our explicit calculations p-wave interlayer tunneling that applies to most of the high-symmetry stacking registries of TMDs as well as InAs/GaSb quantum wells [53, 54, 55]. Different from s-wave tunneling, p-wave interlayer tunneling produces a potential landscape that is second order in the phase angle of the exciton field, leading when no bias voltage is applied to a second-order Josephson effect [56] that breaks the interlayer phase symmetry down to $\mathbb{Z}_{2}$ from $U(1)$. We find that in the large bias limit, the system is described by an effective action in which interlayer $U(1)$ phase symmetry is effectively restored, but p-wave interlayer tunneling leads to a reduced effective band gap and an increased effective temperature for intervalley excitons.
@@
$H_{0}=\sum_{\tau bk}\xi_{bk}a_{\tau bk}^{\dagger}a_{\tau bk}$ (1)
describes the kinetic energy of conduction band electrons and valence band holes. Here $\tau=\pm$ is the valley index and $b=c,v$ is the band (layer) index. $\xi_{ck}=k^{2}/2m_{e}^{*}+E_{g}/2$ and $\xi_{vk}=-k^{2}/2m_{h}^{*}-E_{g}/2$ describe the dispersion of the conduction and valence bands, where $m_{e}^{*}$ and $m_{h}^{*}$ are the effective masses of electrons and holes and $E_{g}$ is the band gap that can be tuned by a perpendicular electric field produced by the difference between top and bottom gate voltages. The next term in the Hamiltonian
@@
describes interlayer tunneling arising from hybridization of electron and hole wavefunctions in the two layers. A nonzero $t_{\tau k}$ explicitly breaks the $U(1)$ symmetry of the model associated with charge conservation in each layer. The momentum and valley dependence of $t_{\tau k}$ depends on symmetries of the system and is crucial for our upcoming results. For most of the high-symmetry stacking registries of angle-aligned TMD homobilayers, direct tunneling is forbidden by rotational symmetry [50, 51], leading to p-wave interlayer tunneling
$t_{\tau k}=v_{t}(\tau k_{x}+ik_{y}).$ (3)
This form of interlayer tunneling also applies to InAs/GaSb quantum wells [53, 54, 55], in which case $\tau$ is the spin index. In this Letter we focus on p-wave interlayer tunneling and briefly discuss other forms of interlayer tunneling at the end. The Coulomb interaction term
$H_{C}=\frac{1}{2A}\sum_{bb^{\prime}\tau\tau^{\prime}}\sum_{kk^{\prime}q}V_{bb^{\prime}}(q)a_{\tau b,k+q}^{\dagger}a_{\tau^{\prime}b^{\prime},k^{\prime}-q}^{\dagger}a_{\tau^{\prime}b^{\prime}k^{\prime}}a_{\tau bk},$ (4)
where $A$ is the system area, distinguishes intralayer ($b=b^{\prime}$) and interlayer ($b\neq b^{\prime}$) interactions but neglects intervalley scattering due to the large momentum transfer required. Electron-hole exchange interactions [57, 58, 59, 60, 61] are also neglected due to the suppression of current matrix elements between electrons and holes in different layers.
@@
Keldysh action.— We derive a nonequilibrium field theory that describes a biased electron-hole bilayer with interlayer tunneling based on the Keldysh formalism [46, 47, 48, 42], outlining here the procedure to obtain the Keldysh action and presenting the main results, with detailed derivations left for the Supplemental Material. We express the model as a path integral along a closed time path $C$ that starts from the distant past, proceeds to the distant future, and then returns to the starting point. The generating function is
@@
where $\mathcal{T}_{C}$ is the contour ordering operator along $C$, and $\rho_{0}$ is the density matrix of the system in the distant past which we take as the equilibrium distribution of decoupled electron and hole layers: $\rho_{0}=e^{-(H_{0}-\mu_{c}N_{c}-\mu_{v}N_{v})/T}$ where $N_{b}=\sum_{\tau k}a_{\tau bk}^{\dagger}a_{\tau bk}$ is the number of electrons in each layer. For notational convenience Eq. (5) is written for a closed system; coupling to leads is included in the theory as the imaginary (dissipative) part of inverse Green functions as detailed in the Supplemental Material. To derive a theory of excitons, we perform a Hubbard-Stratonovich transformation of interlayer electron-hole interactions and introduce the electron-hole pairing fields $\Delta_{kq}^{\tau\tau^{\prime}}$, where $k$ and $q$ are respectively the relative momentum and center-of-mass momentum of an electron-hole pair, and $\tau,\tau^{\prime}$ are the valley indices of electrons and holes. A nonzero value of $\Delta$ reflects spontaneous breaking of interlayer $U(1)$ symmetry associated with formation of the exciton condensate. The $k$-dependence of the pairing fields is irrelevant to the low-energy physics we discuss and is eliminated by projecting the $\Delta$ fields onto the 1s-exciton basis by defining
@@
$\frac{k^{2}}{2m}\varphi_{k}-\frac{1}{A}\sum_{k^{\prime}}V_{cv}(k-k^{\prime})\varphi_{k^{\prime}}=-E_{b}\varphi_{k}.$ (7)
Here $m=m_{e}^{*}m_{h}^{*}/(m_{e}^{*}+m_{h}^{*})$ is the reduced mass of an electron-hole pair and the exciton binding energy $E_{b}$ is defined as the absolute value of the 1s-exciton energy. The 1s-exciton fields $\Phi$ have two valley indices, one for electrons and the other for holes, and we express them in terms of a four-component spinor $(\Phi^{\mu})$ defined as $\Phi^{\tau\tau^{\prime}}=(\sum_{\mu}\Phi^{\mu}\tau_{\mu}/\sqrt{2})^{\tau\tau^{\prime}}$ where $\tau_{0}$ and $\tau_{1,2,3}$ are the $2\times 2$ identity and Pauli matrices in valley space [15]. In this notation $\Phi^{0},\Phi^{3}$ are intravalley exciton fields and $\Phi^{1},\Phi^{2}$ are intervalley exciton fields. Integrating out the fermion fields, we obtain an effective action in terms of the 1s-exciton fields $\Phi$. Following the convention widely used in the literature on Keldysh field theory [47, 48, 42, 46], we transform the forward ($+$) and backward (--) branches of the $\Phi$ fields into the classical (c) and quantum (q) fields defined as
@@
$S_{0}[\bar{\Phi},\Phi]=\frac{1}{A}\sum_{q}\int\frac{d\omega}{2\pi}\,{\rm Tr}\bigg{[}(\omega-\frac{q^{2}}{2M}-E_{g}+E_{b}+i\gamma)\Phi^{\rm q\dagger}_{q}(\omega)\Phi^{\rm c}_{q}(\omega)+{\rm c.c.}+ig(\omega-\mu_{x})\coth\frac{\omega-\mu_{x}}{2T}\Phi^{\rm q\dagger}_{q}(\omega)\Phi^{\rm q}_{q}(\omega)\bigg{]}$ (10)
which describes free excitons with energy $E_{g}-E_{b}+q^{2}/2M$ (where the exciton mass $M=m_{e}^{*}+m_{h}^{*}$) and chemical potential $\mu_{x}$ at temperature $T$. While the quadratic coefficients take the stated form only in the dilute exciton regime (BEC regime) $|E_{g}-E_{b}-\mu_{x}|\ll E_{b}$ and in the frequency range $|\omega-\mu_{x}|\ll E_{b}$, the overall form of Eq. (10) is general and we expect that our qualitative results apply to a larger parameter regime. The imaginary coefficients $\gamma$ and $g$ describe coupling of excitons to leads. Fluctuation-dissipation theorem implies $\gamma=g(\omega-\mu_{x})$. In the absence of interlayer tunneling, the bias voltage $\mu_{x}$ can be absorbed into $\omega$ and the system is equivalent to an unbiased bilayer with a reduced band gap $E_{g}-E_{b}-\mu_{x}$. Excitons spontaneously form and undergo BEC at low enough temperatures when $E_{g}-E_{b}<\mu_{x}$. Below the transition the exciton fields have semiclassical solutions of the form $\Phi^{\rm c}(t)=|\Phi^{\rm c}|e^{-i\mu_{x}t}$ with amplitude determined by the ratio of quadratic and quartic coefficients of the action.
@@
in which the intervalley exciton fields $\Phi^{1},\Phi^{2}$ at frequency $\omega$ are coupled to those at frequency $-\omega$. Because of this coupling, the bias voltage $\mu_{x}$ cannot be absorbed into $\omega$ and the system is out of equilibrium as shown in the Supplemental Material. Diagrammatic representations of $S_{0}$ and $S_{J}$ are shown in Fig. 2(a,b).
@@
Figure 2: Diagrammatic representations of the effective exciton action (a,b) and an electron-hole scattering process due to interlayer tunneling (c). The solid and dotted curves represent fermion and boson fields respectively, the crosses represent interlayer tunneling, and the wavy line represents Coulomb interaction. (a) and (b) respectively represent the free exciton action $S_{0}$ and Josephson action $S_{J}$.
@@
If interlayer tunneling is s-wave, a first-order Josephson term proportional to $\Phi^{\rm q}(\omega=0)$ exists. Second-order terms of the form $\Phi(-\omega)\Phi(\omega)$ also exist and the coefficients are equal for all valley components. For p-wave interlayer tunneling (3), in contrast, angular momentum conservation implies that first-order Josephson terms vanish and that second-order terms are nonzero only for intervalley exciton fields $\Phi^{1},\Phi^{2}$. The second-order Josephson action (11) produces an energy landscape with explicit dependence on the phase angle $\theta=\arg\Phi$ of the form $E_{J}\sim\cos 2\theta$ that breaks the $U(1)$ phase symmetry down to $\mathbb{Z}_{2}$. The interlayer tunneling current satisfies the second-order Josephson relation $I\sim\sin 2\theta$ [56]. For an unbiased electron-hole bilayer below the BEC transition, the exciton fields are static and the system picks one of the two preferred phase angles that differ by $\pi$ as the ground state implying Ising-type phase transitions of the exciton fields. Because the Josephson action (11) involves only intervalley exciton fields, intervalley excitons are energetically favored over intravalley excitons by pinning the phase at one of the two preferred phase angles, in agreement with mean-field theory results in the context of InAs/GaSb quantum wells [54, 55].
@@
Large bias limit.— In the absence of interlayer tunneling, the phase of the exciton field rotates at a constant frequency $\omega=\mu_{x}$. Interlayer tunneling leads to a potential landscape that explicitly breaks the $U(1)$ phase symmetry. The interplay between the $U(1)$ symmetry breaking term that traps the phase of the condensate and the bias voltage that drives a rotating phase gives rise to interesting nonequilibrium physics that is different from previous work on driven-dissipative condensates [41, 42, 43, 44, 45]. For small bias voltage $\mu_{x}$, the exciton condensate is a static one with its phase trapped at one of the potential minima. Above a threshold bias voltage $\mu_{x}\sim c_{J}$ the condensate becomes a dynamical one with rotating phase. The transition from static to dynamical condensates is schematically shown in Fig. 3. If the bias voltage is much larger than the Josephson energy scale $c_{J}$, the phase-dependent energy landscape is swept rapidly by the rotating fields at approximately constant frequency $\omega\approx\mu_{x}$. Instead of Josephson effects, the Josephson action produces an average effect on the exciton fields and $U(1)$ symmetry is effectively restored.
Figure 3: Schematic phase diagram of a biased electron-hole bilayer with interlayer tunneling. $E_{g}$ is the interlayer band gap, $E_{b}$ is the binding energy of interlayer excitons, and $\mu_{x}$ is the bias voltage. The system is out of equilibrium when $\mu_{x}\neq 0$. The blue region represents the region in which exciton condensation occurs, and the color scale represents the strength of excitonic coherence. As the exciton density increases, the condensate undergoes a BEC-BCS crossover [62, 63] and then a Mott transition [64, 65, 66, 67] to an electron-hole plasma. The hatched area represents the large-bias BEC regime in which our theory applies. Inset: schematic plot of exciton density $n_{x}$ as $\mu_{x}$ increases along the red arrow. The solid and dashed curves are the $n_{x}$$\mu_{x}$ curves with and without interlayer tunneling respectively. $\mu_{0}=E_{g}-E_{b}$ is the threshold bias voltage for injection of excitons in the absence of interlayer tunneling, $\mu_{c}$ and $n_{c}$ are the critical bias voltage and critical density for the occurrence of BKT transition, and $\mu_{0}^{\prime}$, $\mu_{c}^{\prime}$, and $n_{c}^{\prime}$ are the corresponding quantities in the presence of interlayer tunneling.
@@
A finer look at the nonequilibrium effects unveils that, despite the small $E_{b}/E_{g}$ ratio, the nonequilibrium dissipative term $\delta T$ can be as important as the $\varepsilon$ term. To see this, we sketch in the inset of Fig. 3 the density of intervalley excitons $n_{x}$ as a function of the bias voltage $\mu_{x}$. The gap reduction for intervalley exciton discussed above shifts the $n_{x}$$\mu_{x}$ curve to the left by $\delta\mu_{0}=\varepsilon$. Exciton condensation occurs when the temperature is below the Berezinskii-Kosterlitz-Thouless (BKT) transition temperature [70, 71, 14, 72] 11footnotetext: Although Eq. (13) was obtained for equilibrium systems, since $U(1)$ symmetry is effectively restored in the large bias limit, Eq. (13) is applicable but with $T$ and $n_{x}$ modified by nonequilibrium effects (Eq. 12).
$T_{\rm BKT}\approx 1.3\,\frac{n_{x}}{M},$ (13)
with $M$ the exciton mass and $n_{x}$ the exciton density. In other words, the critical exciton density for the occurrence of BKT transition at temperature $T$ is $n_{c}\approx MT/1.3$. Due to the effective temperature increase, the critical density increases by $\delta n_{c}\approx M\delta T/1.3$. Since the $n_{x}$$\mu_{x}$ curve is approximately linear at small exciton densities, with the slope approximately given by [15, 38] the geometric capacitance $C=e^{2}\partial n_{x}/\partial\mu_{x}\approx\epsilon/d$ of the bilayer, the critical bias voltage in the presence of interlayer tunneling decreases by
$\delta\mu_{c}=\delta\mu_{0}-\frac{e^{2}}{C}\delta n_{c}\approx\varepsilon-\frac{Me^{2}d}{1.3\,\epsilon}\delta T,$ (14)
@@
The assumption of p-wave interlayer tunneling (3) is crucial for our results and deserves further discussion. Our theory applies to InAs/GaSb quantum wells and angle-aligned TMD homobilayers with four of the six high-symmetry stackings ($R_{h}^{h},R_{h}^{X},H_{h}^{h}$, and $H_{h}^{X}$ [50, 51]), interlayer tunneling is p-wave and our theory is directly applicable. When interlayer tunneling is s-wave (e.g., TMD homobilayers with $R_{h}^{M}$ or $H_{h}^{M}$ stacking), a nonzero first-order Josephson term ($\propto\Phi(\omega=0)$) exists for intravalley excitons, leading to nonzero static exciton density even before the condensation transition occurs. While the tunneling-induced static excitons are not coupled to the high-frequency exciton fields at quadratic level, electrostatic repulsion between excitons leads to an effective gap increase for excitons in both valleys.
For TMD heterobilayers or homobilayers with a nonzero twist angle, the two layers form a moiré pattern with spatially varying local stacking registry. A proper treatment of general TMD bilayers needs to take account of the momentum shift between conduction and valence bands [73, 49, 74, 52] and is left for future work. Excitonic coherence between shifted bands leads to density wave states that break translational symmetry [55, 75, 76]. In a simple intuitive picture, excitons in a moiré potential are localized near one of the high-symmetry stacking sites [77, 52], and the effects of interlayer tunneling are determined by the local stacking registry.